HETAV PATEL

Grade 11,

Iron emits K base α X-ray of energy 6.4 keV and calcium emits K base α X-ray of energy 3.69 keV. Calculate the times taken by an iron K base α photon and a calcium K base α photon to cross through a distance of 3 km

Iron emits K base α X-ray of energy 6.4 keV and calcium emits K base α X-ray of energy 3.69 keV. Calculate the times taken by an iron K base α photon and a calcium K base α photon to cross through a distance of 3 km

Grade:10

1 Answers

Kevin Nash
9 years ago
Distance = 3 km = 3 * 10^3 m C = 3 * 10^8 m/s T = Dist/Speed = 3 * 10^3/3 *10^8 = 10^-5 sec. ⇒ 10 * 10^-8 sec = 10 μs in both case